First Half - D Night

Observers: JHill, BRothberg (Tucson)
Telescope Operator: SAllanson (LBTO)
AO Support: GTaylor (Tucson)
Telescope Support: <a href="https://wiki.lbto.org/bin/edit/Commissioning/MikeW?topicparent=Commissioning.Com20180603;nowysiwyg=0" rel="nofollow" title="MikeW (this topic does not yet exist; you can create it)"> MikeW </a> (LBTO)
Software Support:
Instrument Support: ESolheid (LBTO)
Engineering Support: DCox

Almanac

Sunset - 02:29 (19:29)
Twilight End - 04:04 (21:04)
Night Center - 07:18 (00:18)

Handover to TSIP ~ 00:15

Moon ~ 73%

Plan for On-sky Checks and Calibrations

(CURRENTLY ON-HOLD, WILL CONFIRM A NO-GO/GO at Sunset)

AGW3 Out of Commission

Adjust guidecam_hotspot_x,_y (BGTechProcedures) [0.3 hour] - We never managed to detect light on WFS, See IT 7282

Measure on-axis pupil position and update GCS and IDL parameters. [0.5 hour] WFScam_hotspot_x,_y Use gcs_wfsc_display and GCSGUI to check.

Iterate guidecam_hotspot and WFScam_hotspot with collimation.

Set Probe center position in AGW3 (BGTechProcedures)[0.5 hour, prefer good seeing]

Skip Pupil Wobble measurement as SX M3 mounting is unchanged, but beware IT 7269. (Commissioning/ComTools#PupilWobble)

Check off-axis guiding and collimation. [0.2 hour]

Check LUCI dithers. [0.2 hour] - longslit and imaging

Check LEFTZEROPOINT wrt/LUCI slit [0.5 hour] - MOS masks <a href="https://wiki.lbto.org/bin/edit/Commissioning/StoneJ?topicparent=Commissioning.Com20180603;nowysiwyg=0" rel="nofollow" title="StoneJ (this topic does not yet exist; you can create it)"> StoneJ </a> or <a href="https://wiki.lbto.org/bin/edit/Commissioning/StoneL?topicparent=Commissioning.Com20180603;nowysiwyg=0" rel="nofollow" title="StoneL (this topic does not yet exist; you can create it)"> StoneL </a>

Check location of rotator center on LUCI1. [0.5 hour]

Collect LUCI1 and LUCI2 transform data [2 hours, requires good seeing] (We are not anticipating data reduction during the night.)

In case of bad seeing (> 1.5"):
  • (D-time) H-band spectral library, seeing >>2" - many mags extinction - LUCI-Bino G210-H (Barry)
  • (D-time) Near-IR Local Merger spectroscopy zJHK G200 - LUCI-Bino
  • (E-time) TMS Active with LBC (John H. and Andrew) (seeing <2", up to 3 mags clouds, target flexible)
  • (E-time) LUCI+MODS Binocular preset tests ( IT 6343) - has not been addressed in over a year, but the IT suggests the fix needs to be tested (John H. and Barry) Target: IC 5117 ~ 10min
  • (E-time) LBC+MODS pseudo-monocular test ( IT 5715) - verify that LBC will dither without streaking in exposures (John H. and Barry) Target: Hoag's Object ~ 10min
  • (E-time) LBC - WRS testing on tech chips
  • (E-time) MODS AGW transforms (blue channel both sides)
  • (E-time) obtain zeropoint measurements at J-band for LUCI-1 and LUCI-2 (seeing limited) for Arcetri (goal to constrain where the flux loss is coming from in the optical path).
D-time programs:
  • IMBH Dwarf Galaxy TOO (1 hour) - Barry
  • Dwarf Galaxy Mergers -(LUCI-2 in case of AGW-3 issues) - 3 targets remain in pilot sample -Barry

Preparations

  • Daytime work to investigate no-flux in AGW3
    • AGW3 wfs camera in front of LUCI1 has zero sensitivity to light (or at least 100000 times less than normal)
    • AGW3 still has the high 6K bias level and has normal looking read noise and bias shape.
    • Restarting of software and power cycling has not helped (both last night and today) .
    • Current hypothesis - electronic failure - Lack of clock voltages would make similar images - need spare power supply and CCD controller

Start

UT 00:55 - taking two LUCI-2 dark test exposures, 2.51sec and 10sec luci2.20180604.0001-0004

UT 02:18 - MODS1data computer froze (verified this on Raritan) - ran through troubleshooting, Steve had to hard restart the computer physically, then brought up all the services, took test biases, seems OK. Will open an IT for this incident for reference (IT 7285)

AO Checkout

UT 01:19 - Steve configuring LUCI-2 for AO checkout

UT 01:25 - taking test exposure as per AO checkout (cube 5 exposures ndit=1, dit=2.51)

UT 01:31 - lucii2.20180604.0005-0014 test exposures

Night

UT 03:06 - AO checkout complete - moving to first target SDSSJ1056+3138

luci2.20180604.0030-0045 central wavelength 1.93 (30min)

luci2.20180604.0046-0061) central wavelength 1.96 (40min)

UT 03:23 - Starting science expsoures

UT 03:34 - WFS reporting ~0.67" FWHM/GUIDER 0.6"

UT 03:50 - Current script Cent_wave = 1.93, which cuts off at rest-frame 2.02 microns (1800seconds), second set will do with lambda_central = 1.96 to get the 2.04 [AlIX] line (1800seconds).

Will need to take new set of cals with 1.96 central wavelength

UT 04:20 - DIMM 0.87" Guider 0.6"

UT 04:35 - DIMM 0.85" Guider 0.6"

UT 04:38 - adding 10more minutes to 1.96central wavelength set

UT 04:50 - finished target (70 min total time 30/40 split)

UT 04:51 Moving to Telluric for SDSSJ1056+3138 HD98989 (need two settings for it)

UT 04:57 - starting telluric spectra luci2.20180604.0062-0067 (0064-0065 1.93 central wavelength, 0066-0067 1.96 central wavelength)

==========================================

UT 05:04 - moving to SDSS1307+5426 Position 1 luci2.20180604.0068-0085 - 30minutes shutter time for this position (good S/N LOTS of emission lines)

UT 05:18 - RTD slow again, moving detx = -0.36" to center full object in slit

UT 05:20 - starting spectroscopic exposures

UT 05:37 - DIMM 0.94" Guider 0.74"/WFS 0.75"

UT 05:58 - finished position 1 on SDSS1307+5426

UT 05:59 - moving to Telluric for SDSS1307+5426 (AG +35 1156) luci2.20180604.0086-0090

UT 06:04 - starting spectroscopy of telluric

============================================

UT 06:09 - Moving to Postion 2 of SDSS1307+5426 luci2.20180604.0091-0112 42.5 minutes shutter time

UT 06:19 - Staring spectroscopic exposures (RTD slow again, verified in slit via laptop DS9/IRAF)

UT 06:53 - very very faint continuum, will keep integrating until handover - DIMM 0.81" guider 0.72"

UT 06:56 - finished 12x150sec exposures, going back to start of script

UT 07:12 - stopping script - putting LUCI-2 into safe mode (1 minute early!) - reconfiguring for MODS-Bino

============================================

UT 12:02 - starting calibrations for LUCI-2 G200 HK CenWave 1.96 0.75" slit

luci2.20180604.0112-0116 flats lights off

0117-0121 flats lights on

0122-0123 arcs lights off - don't use - central wavelength wasn't changed to 1.96

0124-0125 arcs lights off

0126-0127 Argon

0128-0129 Xenon

0130-0131 Neon

UT 12:19 - cals complete putting LUCI-2 into safemode

===============================

Successful observing - great data!
Topic revision: r4 - 04 Jun 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback